OpenStudy (anonymous):

whats the answer If ƒ(x ) = 2x, then ƒ -1(x ) = -2x x x - 2

OpenStudy (er.mohd.amir):

f-1(x)=2x-x=x

OpenStudy (anonymous):

whats is the answer ƒ(x ) = 3x + 1 and ƒ -1 = , then ƒ(2) = 1 7 1/3